Revision of Depreciation

Equipment with a cost of $899,100 has an estimated residual value of $89,100, has an estimated useful life of 27 years, and is depreciated by the straight-line method.

a. Determine the amount of the annual depreciation.
$

b. Determine the book value after 14 full years of use.
$

c. Assuming that at the start of the year 15 the remaining life is estimated to be 20 years and the residual value is estimated to be $93,100, determine the depreciation expense for each of the remaining 20 years.
$

Revision of Depreciation

A building with a cost of $371,250 has an estimated residual value of $74,250, has an estimated useful life of 27 years, and is depreciated by the straight-line method.

a. What is the amount of the annual depreciation?
$

b. What is the book value at the end of the fifteenth year of use?
$

c. If at the start of the sixteenth year it is estimated that the remaining life is 15 years and that the residual value is $11,250, what is the depreciation expense for each of the remaining 15 years?
$
Answer #1

a) Annual depreciation = (899100-89100)/27 = $30000 per year

b) Book value after 14 years use = 899100-(30000*14) = $479100

c) Revision of depreciation = (479100-93100)/20 = $19300 per year

2a) Annual depreciation = (371250-74250/27) = $11000 per year

b) Book value after 15th year = 371250-(11000*15) = $206250

c) Depreciation expense = (206250-11250)/15 = $13000 per year